OpenStudy (ny,ny):

which fraction is between 6/10 and 8/10?

OpenStudy (anonymous):

7/10

OpenStudy (ny,ny):

well yes i know that but my answer choices are 3/4, 3/8, 1/3, and 1/6

OpenStudy (ny,ny):

plz explain

OpenStudy (anonymous):

ok we know that 6/10 = 0.6 and 8/10 = 0.8 then just find which one from the choices is between o.6 and 0.8 3/4 = 0.75 3/8 = 0.167 1/3 = 0.333 1/6 = 0.166 so the answer is 3/4

OpenStudy (ny,ny):

ohh lol that was easy thanks

OpenStudy (anonymous):

you're welcome